Circular No. 8154

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ge IN NGC 6122 Further to IAUC 8147, T. Puckett and A. Langoussis report the dis-

covery of an apparent supernova (mag 17.8) on an unfiltered CCD frame taken on June 21.19 UT (confirmed via CCD on June 22.08 by M. Marcus). SN 2003ge is located at α = 16^{h}20^{m}08^{s}.75, δ = +37^{o}47′53′′.7 (equinox 2000.0), which is 8′′.8 west and 0′′.1 north of the center of NGC 6122, but it is not present on images taken by Puckett on 2002 Sept. 13 (limiting mag ∼ 20.5) or on the Palomar Sky Survey (1993 Apr. 30; limiting mag ∼ 21.1).

η CARINAE

Further to IAUC 8146, M. F. Corcoran reports that, since Mar. 6, daily monitoring PCA observations of η Car with the RXTE show six xray peaks with a quasi-period of ∼ 21 days, the last having occurred on June 15 (∼ 19.9 days after the previous x-ray peak) with a 2–10-keV flux of 2.1 × 10^{−10} erg s^{−1} cm^{−2}. On June 19, the 2–10-keV flux had fallen to 7.2 × 10^{−11} erg s^{−1} cm^{−2}; η Car has not been this faint in x-rays since 2002 Feb. 26 — suggesting continuation of the x-ray eclipse ingress (IAUC 8146). A straight-line fit to the June 15–19 data suggests that the x-ray minimum would be reached on June 21.

XTE J1814-338 M. I. Krauss, A. Dullighan, and D. Chakrabarty, Massachusetts In-

stitute of Technology; M. H. van Kerkwijk, University of Toronto; and C. B. Markwardt, University of Maryland and Goddard Space Flight Center, report on a 9700-s observation of XTE J1814-338 (IAUC 8144) on June 20.1 UT with the Chandra X-Ray Observatory (+ ACIS-S/HETGS). The x-ray source remains active, with a 0.5–10-keV flux of ≈ 15 mCrab. Its position is α = 18^{h}13^{m}39^{s}.03, δ = −33^{o}46′22′′.3 (equinox 2000.0), with a 90percent-confidence radius of 0′′.6. The likely optical counterpart was also observed at this position in images acquired on June 6.4 with the Magellan/Clay 6.5-m telescope in Chile, with magnitudes of B = 17.3 and R = 18.8 (uncertainties 0.1 mag). The star brightened by ∼ 0.2 mag in both bands on June 7.4, and it was undetected in a 1989 Digitized Sky Survey image (limiting mag R > 20). Note that there is a nearby field star of similar brightness (R = 19, B = 20) located just 1′′.7 southeast of the proposed counterpart. Further multiwavelength observations of the pulsar as it fades are strongly encouraged.

2003 June 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green